OpenStudy (anonymous):

Let vectors A =(2,1,−4), B=(−3,0,1), and C=(−1,−1,2) What is the angle θAB between vector A and B?

OpenStudy (anonymous):

Take the dot product between vectors A and B \[cos \theta = \frac{A*B}{|A||B|}\]

OpenStudy (anonymous):

so"|A||B|"mean "|A|"=(2,1,4)|B|=(3,0,1)?

OpenStudy (anonymous):

No you are finding the length of the vector which you find using Pythagorean Theorem so \[|A| = \sqrt{2^2 + 1^2 + 4^2}\]

OpenStudy (anonymous):

ok let me calculus

OpenStudy (anonymous):

i use -10/root of 210 and get -0.7 but it was wrong

OpenStudy (anonymous):

nvm i did it !!! thx

OpenStudy (anonymous):

your welcome